OpenStudy (anonymous):

What is the difference of 23x and 9x?

Nnesha (nnesha):

difference = subtract \[\huge\rm (23x)-(9x)\] variables are same so combine their coefficient example 3u+4u = (3+4)u = 7u

OpenStudy (anonymous):

difference = subtraction. whats 23-9. =14. so your difference is 14x

OpenStudy (anonymous):

Difference means to minus. Therefore, 23x−9x =23x+−9x Combine Like Terms =23x+−9x =(23x+−9x) =14x